OpenStudy (anonymous):

A store pays a 6% commission on the first $500 in sales and 8% on sales over $500. Find the commission on an $800 sale.

OpenStudy (***[isuru]***):

800 = 500 + 300 so this will receive the 500 commission + commission above 500 commission for 500 = (6/100)*500 = 30 commission for above 500 ( there is 300 above 500 in this instance) = (8/100)*300 = 24 so the total commission is = 30 + 24 = 56 hope this will help ya !!

OpenStudy (anonymous):

that helps thank you

OpenStudy (***[isuru]***):

u r welcome!